OpenStudy (anonymous):

Hey I will be posting all question on Limits by AK Sanyal(Maths teacher) any of you who can solve any of them will have my appreciation

OpenStudy (amistre64):

this site is more designed for interactions that promote studying as opposed to simply giving our free answers ... just a public service announcement :)

OpenStudy (amistre64):

but feel free to tickle our fancies with something more stimulating than finding the slope between 2 points :)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}\left( (12/(8-x ^{3}) )-(1/(2-x))\right)\] \[\lim_{x \rightarrow 0} \left(\ln (\cos x+\sin x)/(\sqrt{3+2x}-\sqrt{3-7x}) \right) \] \[\lim_{x \rightarrow 3}\left( (\sqrt{x}-\sqrt{3}+\sqrt{x-3})/\sqrt{x ^{2}-9} \right)\] \[\lim_{x \rightarrow \beta}\left( (1-\cos (ax ^{2}+bx+c))/(\beta-x)^{2} \right)\] where \[\alpha\] and \[\beta\] are roots of \[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

feel free to solve them theese are just 4 out 34 such questions

OpenStudy (amistre64):

for ease of reading: \frac{top}{bottom} formats to:\[\frac{top}{bottom}\]

OpenStudy (amistre64):

\[\lim_{x \to 3}\frac{\sqrt{x}-\sqrt{3}+\sqrt{x-3}}{\sqrt{x ^{2}-9}}\]

OpenStudy (anonymous):

this style is much better thank you but how do you find it in the equations panel

OpenStudy (amistre64):

since this goes to 0/0 we can most likely Lhop it

OpenStudy (amistre64):

\[\frac{\top}{bottom}\] just by typing it in; but apparently it reads top is a special word lol

OpenStudy (amistre64):

you can type it directly into the reply box as well; which is what most of us become used to

OpenStudy (anonymous):

In the first one ... 8-x^3 can be written as (2-x)(4+ x^2 + 2x) Now take LCM. We get [ 12- 4 - x^2 - 2x] / (2-x)( 4+ x^2 + 2x) by factorising we get... [ (x+4) ( 2-x) ] / [ (2-x)(4 + x^2 +2x)] 2-x gets cancelled. we are left with (x+4) . now put x=2. we get the value 6.

OpenStudy (amistre64):

conjugates onthe other hand tend to be your friend whin you got sqrts underneath

OpenStudy (amistre64):

in this case tho I think they might not help hmm are you familiar with using Lhopitals rule?

OpenStudy (anonymous):

me? yes. I am familiar with Lhospital's rule.

OpenStudy (amistre64):

that was more directed at the asker tho, but you know; its fair game lol

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=limit+x+to+3+%28sqrt%28x%29-sqrt%283%29%2Bsqrt%28x-3%29%29%2F%28sqrt%28x%5E2-9%29%29 1/sqrt(6) but i got no idea how they did it; it should be in the show steps part tho

OpenStudy (anonymous):

For last is the answer (2beta + b )^2 / 2beta